The rate of return on Cherry Jalopies, Inc., stock over the last five years was 20 percent, 11 percent, -2 percent, 4 percent, and 10 percent. Over the same period, the return on Straw Construction Company’s stock was 16 percent, 21 percent, -3 percent, 4 percent, and 19 percent.

Calculate the variances and the standard deviations for Cherry and Straw. (Enter variance as a decimal and standard deviation as a percent. Round your variance to 5 decimal places and standard deviation to 2 decimal places. Omit the "%" sign in your response.)

   Cherry    Straw
  Variance
  Standard deviation % %
Answer #1
Cherry Straw
Variance    0.00542    0.00866
Standard Deviation 7.36% 9.31%
Working:
# 1 Calculation of average return:
Return
Year Cherry Straw
1      0.2000      0.1600
2      0.1100      0.2100
3     -0.0200     -0.0300
4      0.0400      0.0400
5      0.1000      0.1900
Total (a)      0.4300      0.5700
Average Return (b=a/5)      0.0860      0.1140
# 2 Calculation of Variance:
Cherry:
Return Average Return
Year a b c=(a-b)^2
1      0.2000      0.0860    0.01300
2      0.1100      0.0860    0.00058
3     -0.0200      0.0860    0.01124
4      0.0400      0.0860    0.00212
5      0.1000      0.0860    0.00020
Total (a)    0.02712
Variance (b=a/5)    0.00542
Straw:
Return Average Return
Year a b c=(a-b)^2
1      0.1600      0.1140    0.00212
2      0.2100      0.1140    0.00922
3     -0.0300      0.1140    0.02074
4      0.0400      0.1140    0.00548
5      0.1900      0.1140    0.00578
Total (a)    0.04332
Variance (b=a/5)    0.00866
# 3 Calculation of standard deviation:
Standard deviation = Variance ^ (1/2)
Standard deviation of:
Cherry = Variance ^ (1/2)
=    0.00542 ^ (1/2)
= 7.36%
Straw = Variance ^ (1/2)
=    0.00866 ^ (1/2)
= 9.31%